OpenStudy (sighn0more):

Can someone please explain to me how to work through this problem? Write the equation y - 2 = 4(x + 5) in standard form. Every time I do it I seem to get y - 4x = 22, which I know is incorrect.

OpenStudy (sighn0more):

I know that I have to move all of the constants to the right side of the equation, and by distributing this is what I have so far: y - 2 = 4x + 20. Sadly, I mess it up after that.

OpenStudy (anonymous):

y-2 = 4(x+5) y-2 = 4x+20 you distribute the to what's in the ( ) y= 4x+22 you move the -2 over to the other side by adding it to the 20 and it is now in standard form :)

OpenStudy (anonymous):

* you distribute the 4. Sorry. keyboard messed up on that part and left out the 4... :P

OpenStudy (sighn0more):

Thank you for your help! I think I understand how to work out this problem now :)

OpenStudy (anonymous):

No problem!